OpenStudy (anonymous):

factor completely 7x^3 - 36x^2 + 5x I have divided the x but now I'm lost... any help

OpenStudy (anonymous):

if you divide the whole thing by x you get a quadratic equation so you can use the quadratic formula. Do you know it?

OpenStudy (anonymous):

yes your a life saver ! :)

OpenStudy (mertsj):

x(7x^2-36x+5) x(7x-1)(x-5)

OpenStudy (anonymous):

\[7x^3-36x^2+5x=x(7x^2-36x+5)=x(7x^2-35x-x+5)\] \[=x(7x(x-5)-(x-5))=x(x-5)(7x-1)\]

OpenStudy (anonymous):

thanks :)